Grade 12Algebra

x|x|=4 x =?

ashu Joshi
12 Years agoGrade 12

1 Answer

Sunil Kumar FP
12 Years ago
x|x|=x^2,x>0
=-x^2 x<0
therefore x^2=4x x>0
x=0,4
-x^2=4x x<0
x=-4,0

thanks and regards
sunil kr
askIItian faculty